# Jamming is a first-order transition with quenched disorder in amorphous materials sheared by cyclic quasistatic deformations

### v1

@@ Abstract
Jamming is an athermal transition between flowing and rigid states in amorphous systems such as granular matter, colloidal suspensions, complex fluids and cells. The jamming transition seems to display mixed aspects of a first-order transition, evidenced by a discontinuity in the coordination number, and a second-order transition, indicated by power-law scalings and diverging lengths. Here we demonstrate that jamming is a first-order transition with quenched disorder in cyclically sheared systems with quasistatic deformations. Particle models are simulated in two and three dimensions, which undergo reversible-irreversible transitions under cyclic shear. The ensemble of configurations, with partially crystallized and fragile states excluded, is generated in the irreversible phase, where the system is stationary and particles are diffusive. Detailed scaling analyses are performed on the distribution of coordination numbers. The fluctuation of the jamming density in finite-sized systems has important consequences on the finite-size effects of various quantities, resulting in a square relationship between disconnected and connected susceptibilities, a key signature of first-order transitions with quenched disorder. This study puts the jamming transition into the category of a broad class of transitions in disordered systems, including the first-order transition in the random-field Ising model, brittle yielding of amorphous materials, and melting of ultra-stable glasses.
@@ Introduction.
(ii) Hyperuniformity and associated diverging length scales. Recent studies reveal the spatial distribution of the single-particle contact number, $Z_{i}$, is hyperuniform at jamming, implying a diverging hyperuniform correlation length, $\xi_{\rm H}\sim\Delta Z^{-\nu}$ Hexner et al. (2018, 2019). The value of exponent $\nu$ appears to depend on the way to extract the correlation length and dimensionality Hexner et al. (2018).
@@
(iii) Gardner glass phase, landscape marginality and associated criticality. The marginality at jamming has been established by an independent approach within the framework of replica symmetry breaking Charbonneau et al. (2014); Parisi et al. (2020); Charbonneau et al. (2015a); Berthier et al. (2016, 2019); Li et al. (2021); Urbani et al. (2023). Unjammed hard sphere glasses undergo a Gardner transition where the free energy landscape becomes fractal and marginal, and the caging susceptibility diverges Berthier et al. (2016). The entire Gardner phase, including the jamming limit, is critical. In other words, the caging correlation length $\xi_{\rm G}\sim(\varphi_{\rm J}-\varphi)^{0}$ remains infinite near jamming. The mean-field replica theory predicts the values of exponents in the weak force and small gap distributions, $\alpha=0.41269$ and $\theta=0.42311$, coinciding with the relationship $\theta=1/\alpha-2$ given by the mechanical marginal stability analysis Charbonneau et al. (2014, 2015b); Wang et al. (2022). The theory provides an additional scaling relationship between the cage size $\Delta$ and the entropic pressure $p$, $\Delta\sim p^{-\kappa}$ with $\kappa=1.41574$. The exponents appear to be independent of the dimensionality $d$ for $d\geq d_{\rm u}$ Charbonneau et al. (2014, 2015b), where $d_{\rm u}=2$ is the conjectured upper critical dimension Goodrich et al. (2012).
@@
(iv) Criticality in shear rheology. The criticality of the jamming transition is suggested by scaling analysis of the rheological data obtained in finite-rate shear simulations of flowing states near $\varphi_{\rm J}$ Olsson and Teitel (2007, 2011); Kawasaki et al. (2015). Combing the power-law divergence of the viscosity, $\eta=(\varphi_{\rm J}-\varphi)^{-\beta}$ and the vanishing of yield stress, $\sigma_{\rm Y}\sim(\varphi-\varphi_{\rm J})^{\Delta}$, a critical scaling function is proposed. A diverging rheological correlation length can be extracted from velocity correlations Olsson and Teitel (2007) or non-affine displacements Heussinger and Barrat (2009), $\xi_{\rm R}\sim(\varphi_{\rm J}-\varphi)^{-\nu}$, with $\nu=0.6-1.1$ Olsson and Teitel (2007, 2011); Heussinger and Barrat (2009); Vagberg et al. (2011). The criticality seems to be at odds with the hyperuniformity discussed in (ii) – in the thermodynamical limit, the fluctuations diverge in the former and vanish in the latter.
(v) Discontinuity in the coordination number. The coordination number $Z$, which is considered as an order parameter of the jamming transition, jumps discontinuously from $Z=0$ below jamming, to $Z\geq Z_{\rm iso}$ above, in quasi-static compression or decompression Makse et al. (2000); O’hern et al. (2003). Apparently, as the signature of a first-order transition, this discontinuity is inconsistent with the viewpoints of a continuous transition described above.
@@
It is clear that various diverging length scales have been suggested throughout the literature. However, none of the lengths in (i-iv) can explain the finite-size scaling behavior of the jamming fraction $F_{\rm J}(\varphi,N)$ O’hern et al. (2003); Bertrand et al. (2016); Baity-Jesi et al. (2017); Jin and Yoshino (2021): the data can be reasonably collapsed by a master curve, $F_{\rm J}(\varphi,N)=\mathcal{F}[(\varphi-\varphi_{\rm J})N^{1/2}]$, valid for both compression and shear jamming in two (2D) and three (3D) dimensions. In this study, we show that this scaling can be fully explained by a first-order transition scenario of the jamming transition with quenched disorder. The form $(\varphi-\varphi_{\rm J})N^{1/2}$ originates from the disorder-induced fluctuation of the jamming density itself in finite-sized systems of $N$ particles, which follows the standard central limit theorem. Thus this finite-size scaling is independent of isostaticity, marginality, criticality and hyperuniformity.
@@
Second, in conventional compression quenching protocols, the generated ensemble depends on the initial conditions  Chaudhuri et al. (2010); Ozawa et al. (2012); Jin and Yoshino (2021) and the basins of attraction Bertrand et al. (2016); Ashwin et al. (2012); Martiniani et al. (2017). Here we instead consider an ensemble prepared by cyclic shear, where the states are sampled by well-controlled dynamics similar to that in thermal systems. Recently, the response of amorphous assemblies of particles to cyclic shear has attracted great interest, due to the presence of a nonequilibrium phase transition, called the reversible-irreversible (RI) transition Pine et al. (2005); Corte et al. (2008); Schreck et al. (2013); Nagasawa et al. (2019); Das et al. (2020). Interestingly, the jamming transition lies in the irreversible regime where particle trajectories are asymptotically diffusive Nagasawa et al. (2019); Das et al. (2020). In this study, we restrict our ensemble within the irreversible phase.
Third, in our ensemble, we carefully exclude partially crystallized and fragile states with $Z<Z_{\rm iso}$ that are sensitive to mechanical perturbations or protocol parameters, and regard them as unjammed states with $Z=0$. If such states are included, the discontinuity in $Z$ diminishes and the jamming transition looks continuous, similar to the results reported in Ref. Heussinger and Barrat (2009) obtained by uniform shear. Thus we expect the impactibility between the criticality viewpoint in (iv) and our first-order picture originating from finite-rate effects. For finite-time scales, the existence of transient states with $Z<Z_{\rm iso}$ can lead to a continuous jamming transition Olsson and Teitel (2007, 2011); Heussinger and Barrat (2009). However, our results suggest that, after a sufficiently long time, any configurations with $Z<Z_{\rm iso}$ would eventually relax to unjammed states with zero energy and inter-particle contacts.
@@
We apply cyclic athermal quasistatic shear (CAQS) to standard models of soft, frictionless particles in 2D and 3D (see Models and Shear protocol in Methods). We present 2D data in the main text and 3D data in the Extended Data. For the 2D model, previous studies report a jamming density (J-point density) $\varphi_{\rm J}\approx 0.842-0.843$ O’Hern et al. (2002); Vagberg et al. (2011). The phase diagram of RI transitions near $\varphi_{\rm J}$ is plotted in Fig. 1a. The RI transition is characterized by the one-cycle displacement averaged over particles Nagasawa et al. (2019),
@@
The above results imply that the configurational space is effectively explored by the dynamical trajectory of the system in the irreversible phase, encouraging us to consider a statistical ensemble generated by shear dynamics. In the main text below, we fix $\gamma_{\rm max}={0.7}$, and vary $\varphi$ systematically in the window of $[0.833,0.849]$. At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ independent configurations are collected to construct the ensemble. In the Extended Data, we present additional results obtained for $\gamma_{\rm max}={1}$. All configurations are typically isotropic since they are collected at $\gamma=0$.
@@
Figure 1d shows the time evolution of the coordination number $Z(t)$ in the irreversible phase, obtained from a typical simulation near $\varphi_{\rm J}$. Rattlers (particles with fewer than $d+1$ contacts) are removed in the computation of $Z$. At first glance, one sees the coexistence of jammed ($Z\approx 2d=4$) and unjammed states ($Z=0$), similar to the coexistence of two ferromagnetic states (positive $m$ and negative $m$) in the time evolution of the magnetization $m(t)$ in an Ising model near a first-order phase transition Binder (1987a). However, a more careful examination reveals four distinct states, which we discuss below.
Figure 1: Cyclic athermal quasi-static shear simulations. (a) Reversible ($\delta r_{\infty}<0.1$, blue region) and irreversible ($\delta r_{\infty}>0.1$, yellow region) phases ($N={1000}$). The solid line represents $\varphi_{\rm J}^{\infty}={0.8432}$, and the dashed line represents $\gamma_{\rm max}=0.7$ used to generate the ensemble in the main text. (b) One-cycle displacement $\delta r(t)$ and (d) average coordination number $Z(t)$ for a typical sample. (c) The MSD data show diffusive behavior $\langle\Delta r^{2}(t)\rangle\sim t$ (line). Data in (b-c) are obtained at a fixed $\varphi=0.841$.
@@
Figure 2: Four states. (a) Semi-log plot of the probability distribution $p(Z)$. The four peaks of unjammed (U), partially crystallized (C), fragile (F) and jammed (J) states, are indicated. The inset shows a linear plot of $p(Z)$ without the unjammed peak. Fractions are plotted as a function of $\varphi$ in (b), and of $t$ in (c). In (c), each data point is obtained for a small window $[t-\delta t,t+\delta t]$ with $\delta t={250}$. (d) Peak coordination numbers $Z_{\rm J}^{*}$ of jammed states (upper branch, open symbols) and $Z_{\rm F}^{*}$ of fragile states (lower branch, filled symbols) as functions of $\varphi$, for $N=256$ (diamonds), 512 (triangles), 1000 (squares), and 2000 (circles). In (e-h), we show typical configurations of (e) partially crystallized, (f) fragile percolating in one direction, (g) fragile percolating in both directions, and (h) jammed states. Contact forces are represented by bonds, whose width is proportional to the magnitude of force. The red and blue disks are non-rattler and rattler particles respectively. We set $\varphi=0.841$ for (a) and (c), $N=256$ for (b) and (c), and $N=1000$ for (e-h).
@@
(iv) Fragile states. We define the states in the broad peak $43<Z<$ 4, excluding the crystalline peak $p_{\rm C}(Z)$ as the fragile states.
@@
In Fig. 2d, we plot the maximum coordination numbers, $Z_{\rm J}^{*}$ and $Z_{\rm U}^{*}$, of jammed and fragile peaks. Around $\varphi_{\rm J}$, there is a small gap $\delta Z_{\rm gap}\approx 0.2$ between $Z_{\rm J}^{*}$ (upper branch) and $Z_{\rm U}^{*}$ (lower branch). The results are very similar to those obtained by quasistatic uniform shear in Ref. Heussinger and Barrat (2009) (see Fig. 4 therein). In Ref. Heussinger and Barrat (2009), it is suggested that $\delta Z_{\rm gap}$ vanishes in the thermodynamic limit, and thus the jamming transition is continuous. However, as we show below, the fragile states are generated due to incomplete energy minimization and are mechanically unstable. Once such fragile states are excluded, the lower branch only contains unjammed states with $Z_{\rm U}^{*}=0$, which is separated by a large gap $\delta Z_{\rm gap}\approx 4$ from the upper branch $Z_{\rm J}^{*}\geq 4$.
@@
Previously, the fragile states were obtained by uniform shear at non-zero strains $\gamma>0$ below $\varphi_{\rm J}$ in experiments Bi et al. (2011) and simulations Kumar and Luding (2016). It was proposed that fragile and jammed states differ in the percolation of the strong force network: in fragile states the percolation occurs anisotropically in one direction only, while in jammed states it occurs isotropically in both directions. In this study, all states are collected during cyclic shear at $\gamma=0$, without systematically introducing anisotropy. Indeed, the fragile states with $43<Z<$ 4 can have percolated force networks of non-rattlers in one or two directions (see Fig. 2f and g). Thus in our case, anisotropy cannot effectively distinguish fragile from jammed states.

### v2

@@ Abstract
Jamming is an athermal transition between flowing and rigid states in amorphous systems such as granular matter, colloidal suspensions, complex fluids and cells. The jamming transition seems to display mixed aspects of a first-order transition, evidenced by a discontinuity in the coordination number, and a second-order transition, indicated by power-law scalings and diverging lengths. Here we demonstrate that jamming is a first-order transition with quenched disorder in cyclically sheared systems with quasistatic deformations, in two and three dimensions. Based on scaling analyses, we show that fluctuations of the jamming density in finite-sized systems have important consequences on the finite-size effects of various quantities, resulting in a square relationship between disconnected and connected susceptibilities, a key signature of the first-order transition with quenched disorder. This study puts the jamming transition into the category of a broad class of transitions in disordered systems where sample-to-sample fluctuations dominate over thermal fluctuations, suggesting that the nature and behavior of the jamming transition can be understood within the developed theoretical framework of the athermally driven random-field Ising model.
@@ Introduction.
(ii) Hyperuniformity and associated diverging length scales. Recent studies reveal that the spatial distribution of the single-particle contact number, $Z_{i}$, is hyperuniform at jamming. The contact hyperuniformity is established by two power-law scalings measured in simulated packings Hexner et al. (2018, 2019): the scaling of the fluctuations of the average contact number in a hyper-cube of volume $\ell^{d}$, $\sigma_{Z}(\ell)\sim\ell^{-\mu}$ with $\mu=1$, and the small wave-vector scaling of the contact number structure factor, $S_{Z}(q)\sim q^{\alpha_{Z}}$ with $\alpha_{Z}\approx 1.53$. The crossover from the hyperuniform regime ($\ell\ll\xi_{\rm H}$) to the uniform regime ($\ell\gg\xi_{\rm H}$) defines a hyperuniform correlation length $\xi_{\rm H}$, which diverges at the jamming transition, $\xi_{\rm H}\sim\Delta Z^{-\nu_{\rm H}}$. The value of the exponent $\nu_{\rm H}$ appears to depend on how to extract the correlation length and the dimensionality.
@@
(iii) Gardner glass phase, landscape marginality and associated criticality. The marginality at jamming has been established by an independent approach within the framework of replica symmetry breaking Charbonneau et al. (2014); Parisi et al. (2020); Charbonneau et al. (2015a); Berthier et al. (2016, 2019); Li et al. (2021); Urbani et al. (2023). Unjammed hard sphere glasses undergo a Gardner transition where the free energy landscape becomes fractal and marginal, and the caging susceptibility diverges Berthier et al. (2016). The entire Gardner phase, including the jamming limit, is critical. In other words, the caging correlation length, $\xi_{\rm G}\sim\infty$, remains infinite near jamming, when the transition is approached ($\varphi\rightarrow\varphi_{\rm J}$) from below ($\varphi<\varphi_{\rm J}$). The mean-field replica theory predicts the values of exponents in the weak force and small gap distributions, $\alpha=0.41269$ and $\theta=0.42311$, coinciding with the relationship $\theta=1/\alpha-2$ given by the mechanical marginal stability analysis Charbonneau et al. (2014, 2015b); Wang et al. (2022). The theory provides an additional scaling relationship between the cage size $\Delta$ and the entropic pressure $p$, $\Delta\sim p^{-\kappa}$ with $\kappa=1.41574$. The exponents appear to be independent of the dimensionality $d$ for $d\geq d_{\rm u}$ Charbonneau et al. (2014, 2015b); Parisi et al. (2018), where $d_{\rm u}=2$ is the conjectured upper critical dimension of the jamming transition Goodrich et al. (2012).
@@
(iv) Criticality in shear rheology. The criticality of the jamming transition is suggested by scaling analyses of the rheological data obtained in finite-rate shear simulations of flowing states near $\varphi_{\rm J}$ Olsson and Teitel (2007, 2011); Kawasaki et al. (2015). Combing the power-law divergence of the viscosity, $\eta=(\varphi_{\rm J}-\varphi)^{-\beta}$ and the vanishing of yield stress, $\sigma_{\rm Y}\sim(\varphi-\varphi_{\rm J})^{\Delta}$, a critical scaling function is proposed. A diverging rheological correlation length can be extracted from velocity correlations Olsson and Teitel (2007) or non-affine displacements Heussinger and Barrat (2009), $\xi_{\rm R}\sim(\varphi_{\rm J}-\varphi)^{-\nu}$, with $\nu\approx 1$  Olsson and Teitel (2007); Vagberg et al. (2011); Olsson and Teitel (2020). The criticality seems to be at odds with the hyperuniformity discussed in (ii) – in the thermodynamical limit, the fluctuations diverge in the former and vanish in the latter.
(v) Discontinuity in the coordination number. The coordination number $Z$, which is considered as an order parameter of the jamming transition Liu and Nagel (2010), jumps discontinuously from $Z=0$ below jamming, to $Z\geq Z_{\rm iso}$ above, under quasi-static compression/decompression Makse et al. (2000); O’hern et al. (2003) or shear Vinutha et al. (2020). In unjammed states, $Z=0$ because particles can push each other apart, leaving no overlapping between them. In jammed states, the contact network spans the entire system and the minimum condition to have such networks is $Z=Z_{\rm iso}$. Thus the discontinuous jump in $Z$ at jamming is essentially related to isostaticity. Apparently, as the signature of a first-order transition, this discontinuity is inconsistent with the viewpoint of a continuous transition described above. Note that not all physical quantities exhibit a discontinuity at jamming: for example, the pressure $P$ and the shear modulus $G$ vanish continuously when the jamming transition is approached from above Makse et al. (2000); O’hern et al. (2003). On the other hand, other quantities, such as the bulk modulus $B$ and the fraction of non-rattlers $f_{\rm NR}$, jump abruptly at jamming, similar to $Z$.
@@
It is clear that various diverging length scales have been suggested throughout the literature. However, none of the lengths in (i-iv) can explain the finite-size scaling behavior of the jamming fraction $F_{\rm J}(\varphi,N)$ O’hern et al. (2003); Bertrand et al. (2016); Baity-Jesi et al. (2017); Jin and Yoshino (2021): the data can be reasonably collapsed by a scaling form, $F_{\rm J}(\varphi,N)=\mathcal{F}[(\varphi-\varphi_{\rm J})N^{1/2}]$, valid for both compression and shear jamming in two (2D) and three (3D) dimensions. In this study, we show that this scaling can be fully explained by a first-order transition scenario of the jamming transition with quenched disorder. The form $(\varphi-\varphi_{\rm J})N^{1/2}$ originates from the disorder-induced fluctuation of the jamming density itself in finite-sized systems of $N$ particles, which follows the standard central limit theorem. Thus this finite-size scaling is independent of isostaticity, marginality, criticality and hyperuniformity.
@@
Second, in conventional compression protocols, the generated ensemble depends on the initial conditions  Chaudhuri et al. (2010); Ozawa et al. (2012); Jin and Yoshino (2021) and the basins of attraction Bertrand et al. (2016); Ashwin et al. (2012); Martiniani et al. (2017). Here we instead consider an ensemble prepared by cyclic shear, where the states are sampled by well-controlled dynamics similar to those in thermal systems. Recently, the response of amorphous assemblies of particles to cyclic shear has attracted great interest, due to the presence of a nonequilibrium phase transition, called the reversible-irreversible (RI) transition Pine et al. (2005); Corte et al. (2008); Schreck et al. (2013); Nagasawa et al. (2019); Das et al. (2020). Interestingly, the jamming transition lies in the irreversible regime where particle trajectories are asymptotically diffusive Nagasawa et al. (2019); Das et al. (2020). In this study, we restrict our ensemble within the irreversible phase. The jamming density of this ensemble is equivalent to the minimum jamming density, or the jamming-point (J-point) density that is obtained by rapid quench O’hern et al. (2003).
Third, in our ensemble, we carefully exclude partially crystallized and fragile states with $Z<Z_{\rm iso}$ that are sensitive to mechanical perturbations or protocol parameters, and regard them as unjammed states with $Z=0$. If such states are included, the discontinuity in $Z$ diminishes and the jamming transition looks continuous, similar to the results reported in Ref. Heussinger and Barrat (2009) obtained by uniform shear. Thus we expect the impactibility between the criticality viewpoint in (iv) and our first-order picture originating from finite-rate effects. For finite-time scales, the existence of transient states with $Z<Z_{\rm iso}$ can lead to a continuous jamming transition Olsson and Teitel (2007, 2011); Heussinger and Barrat (2009). However, our results suggest that, after a sufficiently long time, any configurations with $Z<Z_{\rm iso}$ would eventually relax to unjammed states with zero energy and inter-particle contacts. Note that the timescale for the system to attain force balance diverges at the shear jamming transition Vinutha et al. (2020), which means that the real quasistatic limit would correspond to extremely small shear rates in large systems.
Figure 1: Cyclic athermal quasi-static shear simulations. (a) Reversible ($\delta r_{\infty}<0.1$, blue region) and irreversible ($\delta r_{\infty}>0.1$, yellow region) phases ($N={1000}$). The solid line represents $\varphi_{\rm J}^{\infty}={0.8432}$, and the dashed line represents $\gamma_{\rm max}=0.7$ used to generate the ensemble in the main text. (b) One-cycle displacement $\delta r(t)$ and (d) coordination number $Z(t)$ for a typical sample, at $\varphi=0.841$. (c) Several typical particle trajectories during 100 cycles at $\varphi=0.8425$. Positions are recorded at $\gamma=0$ only, and the system size is $43.8\times 43.8$ with periodic boundary conditions. (inset) The MSD data show diffusive behavior $\langle\Delta r^{2}(t)\rangle\sim t$ (line).
@@
We apply cyclic athermal quasistatic shear (CAQS) to standard models of soft, frictionless particles in 2D and 3D (see Models and Shear protocol in Methods). We present 2D data in the main text and 3D data in the Supplementary Information (SI). For the 2D model, previous studies report a jamming density (J-point density) $\varphi_{\rm J}\approx 0.842-0.843$ O’Hern et al. (2002); Vagberg et al. (2011). The phase diagram of RI transitions near $\varphi_{\rm J}$ is plotted in Fig. 1a. The RI transition is characterized by the one-cycle displacement averaged over particles Nagasawa et al. (2019),
@@
The above results imply that the configurational space is effectively explored by the dynamical trajectory of the system in the irreversible phase, encouraging us to consider a statistical ensemble generated by shear dynamics. In this study, we fix $\gamma_{\rm max}={0.7}$ unless otherwise specified, and vary $\varphi$ systematically in the window of $[0.833,0.849]$. At each $\varphi$, in total $t_{\rm max}\times N_{\rm s}$ independent configurations are collected to construct the ensemble. In the SI, we present additional results obtained for $\gamma_{\rm max}={1}$. All configurations are typically isotropic since they are collected at $\gamma=0$.
@@
Figure 1d shows the evolution of the coordination number $Z(t)$ in the irreversible phase, obtained from a typical simulation near $\varphi_{\rm J}$. Rattlers (particles with fewer than $d+1$ contacts) are removed in the computation of $Z$. At first glance, one sees the coexistence of jammed ($Z\approx 2d=4$) and unjammed states ($Z=0$), similar to the coexistence of two ferromagnetic states (positive $m$ and negative $m$) in the time evolution of the magnetization $m(t)$ in an Ising model near a first-order phase transition Binder (1987a). However, a more careful examination reveals four distinct states, which we discuss in the next section.
Figure 2: First three shear cycles. The (a) strain $\gamma$, (b) sample-averaged coordination number $Z$, (c) sample-averaged pressure $P$, and (d) jamming fraction $F_{\rm J}$ are plotted as functions of $t$ during the first three shear cycles ($\varphi=0.8425$, $N_{\rm s}=768$ samples). Orange and blue curves are obtained from rapidly quenched ($\varphi_{\rm j}=\varphi_{\rm J}\approx 0.843$) and mechanically trained ($\varphi_{\rm j}\approx 0.848$) initial configurations respectively. The jamming fraction $F_{\rm J}$ used in the scaling analysis is obtained at $\gamma=0$ (black cycles in (d)). In (e) and (f), plotted is the $Z(t)$ of two typical samples initially generated by rapid quenching. The system status at $\gamma=0$ is highlighted by black cycles, which fluctuate among jammed ($Z\geq 4$), fragile ($43<Z<$ 4) and unjammed ($Z=0$) states.
It is well known now that the jamming density $\varphi_{\rm j}$ is not unique and protocol-dependent Parisi and Zamponi (2010); Chaudhuri et al. (2010); Ozawa et al. (2012); Kumar and Luding (2016); Kawasaki and Miyazaki (2020); Das et al. (2020); Babu et al. (2021); Jin and Yoshino (2021) (In this study, $\varphi_{\rm J}$ denotes the J-point density that is the minimum jamming density, and $\varphi_{\rm j}$ denotes the protocol-dependent jamming density): In the compression protocol, $\varphi_{\rm j}$ depends on the compression rate or the density of the initial equilibrium configuration Chaudhuri et al. (2010); Ozawa et al. (2012); in the cyclic athermal quasistatic compression (CAQC) protocol, $\varphi_{\rm j}$ depends on the volumetric strain amplitude (the maximum density $\varphi_{\rm max}$ to which the system is compressed) and the number $N_{\rm c}$ of cycles Kumar and Luding (2016); Kawasaki and Miyazaki (2020); in the CAQS protocol, $\varphi_{\rm j}$ depends on the shear strain amplitude $\gamma_{\rm max}$ and $N_{\rm c}$ Das et al. (2020); Babu et al. (2021). The continuous range of possible $\varphi_{\rm j}$ is called a jamming-line (J-line), and the minimum jamming density on the J-line is the J-point density $\varphi_{\rm J}$ realized by rapid quenching Babu et al. (2021). The protocol-dependence of $\varphi_{\rm j}$ in the current model has been investigated in Ref. Kawasaki and Miyazaki (2020) using the CAQC protocol, and the maximum jamming density obtained there is $\varphi_{\rm j}\approx 0.8465$ after one over-compression cycle, compared to $\varphi_{\rm J}\approx 0.842-0.843$. In order to examine the effects of non-universal $\varphi_{\rm j}$, in this study we prepare two types of samples: rapidly quenched samples with $\varphi_{\rm j}=\varphi_{\rm J}\approx 0.843$, and mechanically trained samples generated by the CAQC protocol with $\varphi_{\rm j}\approx 0.848>\varphi_{\rm J}$.
The protocol-dependence of $\varphi_{\rm j}$ is called a memory effect in Ref. Kumar and Luding (2016). Below we demonstrate that the cyclic shear simulation employed in this study erases the memory of the initial state: given the initial $\varphi_{\rm j}\approx 0.848>\varphi_{\rm J}$, the jamming density reduces to $\varphi_{\rm J}\approx 0.843$ after one complete shear cycle. Figure 2b-d shows the evolution of the sample-averaged coordination number $Z$, the sample-averaged pressure $P$, and the jamming fraction $F_{\rm J}$ of all considered samples at the given $t$. The differences between the curves of rapidly quenched ($\varphi_{\rm j}=\varphi_{\rm J}\approx 0.843$) and mechanically trained systems ($\varphi_{\rm j}\approx 0.848$) only appear in the first cycle. Thus the effect of the variable $\varphi_{\rm j}$ is eliminated during the construction of the ensemble by cyclic shear. Previous studies have shown that, independent of the initial jamming density $\varphi_{\rm j}$, if the system is sheared by large strain deformations beyond the yield strain $\gamma_{\rm Y}$, it always evolves into a critical steady state that has a generic jamming density at $\varphi_{\rm J}$, and such a process is accompanied with significant shear dilatancy or hardening effects Babu et al. (2021); Pan et al. (2023b); Xing et al. (2024). According to Fig. 2b-d, the system has indeed reached the steady state before the strain reversal, as the strain amplitude $\gamma_{\rm max}=0.7$ is larger than the yield strain $\gamma_{\rm Y}\sim 0.1$.
Next we illustrate that the statistical properties of the ensemble are determined by fluctuation effects. The overall behavior of the sample-averaged $Z(t)$ and $P(t)$ is very similar to that reported in previous cyclic shear simulations Kumar and Luding (2016); Luding et al. (2021): one observes periodic patterns with the sharp onset of unjamming upon strain reversal at integer or half-integer $t$. However, no regular patterns can be identified in single-sample curves, which look more like stochastic processes (see Fig. 2e-f). Similar behavior has been observed in previous simple shear simulations Heussinger and Barrat (2009) and experiments Behringer et al. (2008), showing that the coexistence of jammed and unjammed states during shear is a generic property of the system, at a fixed $\varphi$ near the jamming transition. It is useful to consider two types of fluctuations. (i) Fluctuations between different cycles in the simulation of a given sample. The configurations at consecutive strain steps $t$ and $t+\delta t$ are related by both affine and non-affine deformations. Due to non-affine deformations, the two finite-size configurations can have different structures and jamming densities $\varphi_{\rm J}^{N}$. If the preset constant density $\varphi$ in the CAQS simulation is close to the average $\varphi_{\rm J}$, then the two consecutive configurations can be either jammed ($\varphi_{\rm J}^{N}<\varphi$) or unjammed ($\varphi_{\rm J}^{N}>\varphi$). This fluctuation effect is the origin of stochastic-like behavior in Fig. 2e-f. (ii) Sample-to-sample fluctuations at a given $t$. For similar reasons, at a given $t$, two individual samples can be either jammed or unjammed due to sample-to-sample fluctuations (see marked points at the integer $t$ in Fig. 2e-f). The jamming fraction $F_{\rm J}$ due to such sample-to-sample fluctuations is plotted in Fig. 2d, whose scaling behavior is analyzed in detail below. Note that the fluctuations in (i) and (ii) essentially have the same origin: in finite-size systems, the jamming density $\varphi_{\rm J}^{N}$ has a distribution $p(\varphi_{\rm J}^{N})$ due to the existence of abundant amorphous states.
As shown in Fig. 2d, upon on the strain reversal, $F_{\rm J}$ rapidly decays to zero, which means that the configuration jammed along a given shear direction is unstable to reverse shear deformations Kumar and Luding (2016); Boschan et al. (2019); Seto et al. (2019). This reversal-unjamming effect occurs within a strain interval $\delta\gamma\lesssim 0.2$, and after this interval $F_{\rm J}$ reaches a stationary plateau. Because the maximum strain $\gamma_{\rm max}>0.2$ in our protocol, the configurations collected in the ensemble are always in the stationary regime.
@@
Figure 3: Four states. (a) Semi-log plot of the probability distribution $p(Z)$. The four peaks of unjammed (U), partially crystallized (C), fragile (F) and jammed (J) states, are indicated. The inset shows a linear plot of $p(Z)$ without the unjammed peak. Fractions are plotted as a function of $\varphi$ in (b), and of $t$ in (c). In (c), each data point is obtained for a small window $[t-\delta t,t+\delta t]$ with $\delta t={250}$. (d) Peak coordination numbers $Z_{\rm J}^{*}$ of jammed states (upper branch, open symbols) and $Z_{\rm F}^{*}$ of fragile states (lower branch, filled symbols) as functions of $\varphi$, for $N=256$ (diamonds), 512 (triangles), 1000 (squares), and 2000 (circles). In (e-h), we show typical configurations of (e) partially crystallized, (f) fragile percolating in one direction, (g) fragile percolating in both directions, and (h) jammed states. Contact forces are represented by bonds, whose width is proportional to the magnitude of force. The red and blue disks are non-rattler and rattler particles respectively. We set $\varphi=0.841$ for (a) and (c), $N=256$ for (b) and (c), and $N=1000$ for (a, e-h).
@@
(iv) Fragile states. We define the states in the broad peak $43<Z<$ 4, excluding the crystalline peak $p_{\rm C}(Z)$, as the fragile states.
@@
In Fig. 3d, we plot the maximum coordination numbers, $Z_{\rm J}^{*}$ and $Z_{\rm F}^{*}$, of jammed and fragile peaks. Around $\varphi_{\rm J}$, there is a small gap $\delta Z_{\rm gap}\approx 0.2$ between $Z_{\rm J}^{*}$ (upper branch) and $Z_{\rm F}^{*}$ (lower branch). The results are very similar to those obtained by quasistatic uniform shear in Ref. Heussinger and Barrat (2009) (see Fig. 4 therein). In Ref. Heussinger and Barrat (2009), it is suggested that $\delta Z_{\rm gap}$ vanishes in the thermodynamic limit, and thus the jamming transition is continuous. However, as we show below, the fragile states are generated due to incomplete energy minimization and are mechanically unstable. Once such fragile states are excluded, the lower branch only contains unjammed states with $Z_{\rm U}^{*}=0$, which is separated by a large gap $\delta Z_{\rm gap}\approx 4$ from the upper branch $Z_{\rm J}^{*}\geq 4$.
@@
Previously, the fragile states were obtained by uniform shear at non-zero strains $\gamma>0$ below $\varphi_{\rm J}$ in experiments Bi et al. (2011) and simulations Kumar and Luding (2016). It was proposed that fragile and jammed states differ in the percolation of the strong force network: in fragile states the percolation occurs anisotropically in one direction only, while in jammed states it occurs isotropically in both directions. In this study, all states are collected during cyclic shear at $\gamma=0$, without systematically introducing anisotropy. Indeed, the fragile states with $43<Z<$ 4 can have percolated force networks of non-rattlers in one or two directions (see Fig. 3f and g). Thus in our case, anisotropy cannot effectively distinguish fragile from jammed states. It should be noted that, due to strong fluctuation effects, the fragile states can be generated either before or after jamming during the shear procedure (see Fig. 2e-f).